OpenStudy (anonymous):

Three noncollinear points determine a triangle. How many triangles can be formed with 8 points, no three of which are collinear? A. 56 B. 24 C. 336 D. 6720

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 8 C 3 = (8!)/(3!(8-3)!) 8 C 3 = (8!)/(3!5!) 8 C 3 = (8*7*6*5!)/(3!5!) 8 C 3 = (8*7*6)/(3!) 8 C 3 = (8*7*6)/(3*2*1) 8 C 3 = (8*7*2)/(2*1) 8 C 3 = (8*7)/(1) 8 C 3 = 56 So there are 56 different triangles.